C3H8 combusts. molar masses:
tricarbon octahydride: 44.11 g/moloxygen gas: 32 g/molwater: 18.02 g/molcarbon dioxide: 44.01 g/mol
a. Write a balanced chemical reaction.
b. If 5.34 g of tricarbon octahydride reacts with 25.2 g of oxygen gas, what is the limiting reactant? (tricarbon octahydride)
c. How many grams of water is produced? (8.72168)
d. How much excess reactant is left over after the reaction is complete? (5.84)
e. What is the percent yield of water if 6.98 g of water is actually produced in a laboratory experiment? (80.0459)

1
Expert's answer
2021-03-26T05:44:57-0400

Solution:

(a):

The balanced chemical equation:

C3H8(g) + 5O2(g) → 3CO2(g) + 4H2O(l)


(b):

Determine moles of each reactant:

(5.34 g C3H8) × (1 mol C3H8 / 44.11 g C3H8) = 0.1211 mol C3H8

(25.2 g O2) × (1 mol O2 / 32.00 g O2) = 0.7875 mol O2


According to the chemical equation above: n(C3H8) = n(O2)/5

Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose C3H8:

n(O2) = 5 × n(C3H8) = 5 × 0.1211 mol = 0.6055 mol


The calculation above means that we need 0.6055 mol of O2 to completely react with 0.1211 mol C3H8.

We have 0.7875 mol O2 and therefore more than enough oxygen.

Thus oxygen (O2) is in excess and tricarbon octahydride (C3H8) must be the limiting reactant.

The limiting reactant is tricarbon octahydride (C3H8).


(c):

(5.34 g C3H8) × (1 mol C3H8 / 44.11 g C3H8) × (4 mol H2O/ 1 mol C3H8) × (18.02 g H2O / 1 mol H2O) = 8.726 g H2O

8.726 grams of water (H2O) is produced.


(d):

0.7875 mol O2 - 0.6055 mol of O2 = 0.182 mol O2 (excess O2)

(0.182 mol O2) × (1 mol O2 / 32.00 g O2) = 5.824 g O2

5.824 grams of oxygen gas (O2) is left over after the reaction is complete.

(e):

%H2O = (6.98 g / 8.726 g) × 100% = 79.99% = 80.00%

The percent yield of water (H2O) is 80.00%.